# Unlocking State-Tracking in Linear RNNs Through Negative Eigenvalues

### v2

@@ Acknowledgments
We would like to thank David Salinas, Herilalaina Rakotoarison, Eric Alcaide, Arya Akhavan, Matia Bojovic, Erfan Mirzaei and the active members of the Flash Linear Attention discord channel for their constructive discussions and feedback. We also wish to thank Mattia Pini for helping us with the Franklin computing cluster at IIT, on which we run our largest experiments. This research was partially supported by the following sources: PNRR MUR Project PE000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)“, funded by the European Union – NextGenerationEU, and EU Project ELSA under grant agreement No. 101070617. TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Frank Hutter acknowledges financial support by the Hector Foundation. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. The authors gratefully acknowledge the Gauss Center for Supercomputing eV (www.gauss-centre.eu) for funding this project by providing computing time on the GCS supercomputer JUWELS at Jülich Supercomputing Center (JSC). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
@@
  • Appendix A contains additional details on the notation used, on Table 1, the relationship between RNNs and regular languages, the assumption of finite precision, the states, and the decoder function.
@@ A.2 Details of Table 1
The Mamba recurrence in Equations 3 and 4 in (Gu & Dao, 2023) is applied independently to each channel of the input sequence. Expressing the full recurrence in the matrix-form of (1) is challenging, as it would require concatenating the rows of the matrix ${\bm{H}}_{t}$. For simplicity, in Table 1 we write instead the recurrence for each row of ${\bm{H}}_{t}$. In particular, Let ${\bm{x}}_{t}\in\mathbb{R}^{d}$ be the input of the layer, ${\bm{W}}_{\Delta}\in\mathbb{R}^{d\times d}$, ${\bm{w}}_{2}\in\mathbb{R}^{d}$, ${\bm{W}}_{1}=({\bm{w}}_{1},\dots,{\bm{w}}_{n})^{\top}\in\mathbb{R}^{n\times d}$ be learnable parameters, ${\bm{q}}_{t}\in\mathbb{R}^{n},{\bm{k}}_{t}=(k_{t,1},\dots,k_{t,n})^{\top}\in\mathbb{R}^{n}$ be learnable functions of the input and $\bm{\Delta}_{t}=\mathrm{softplus}({\bm{W}}_{\Delta}{\bm{x}}_{t})$. Then, if we set ${\bm{H}}_{t}=({\bm{h}}_{t,1},\dots,{\bm{h}}_{t,n})^{\top}\in\mathbb{R}^{n\times d}$ and ${\bm{H}}_{0}=0$, we can write the recurrence for the $i$-th row of ${\bm{H}}_{t}$ and the output as
@@
where $\mathbf{1}$ is the vector of $n$ ones. However, such a recurrence is not in the form (1), since we have replaced the matrix-matrix product ${\bm{A}}({\bm{x}}_{t}){\bm{H}}_{t}$ with the element-wise product ${\bm{A}}({\bm{x}}_{t})\odot{\bm{H}}_{t}$. Note that we follow the implementation of ${\bm{B}}({\bm{x}}_{t})$ used in the official Mamba codebase, which simplifies the expression originally presented in Equation 4 of (Gu & Dao, 2023) as described by the authors in a GitHub Issue [^3] https://github.com/state-spaces/mamba/issues/19.
@@ A.3 Regular Languages and Recurrent Neural Networks
It is apparent that any FSA can be implemented by an RNN layer if $g$ is sufficiently expressive to model its state transition function.
@@ A.4.1 Initial State, Matrix-valued States, and The Decoder function
While we mention that the states ${\bm{H}}_{t}$ are generally matrices of dimension $n\times d$, for our theoretical constructions (excluding the first two theorems), we set $d=1$, so that states are vector-valued. Hence, for the problems that we consider, we find that having a matrix-valued state ($d>1$) brings no theoretical advantage.
To compute the output $\hat{\bm{y}}_{t}$ from the state ${\bm{H}}_{t}$ and the vector ${\bm{x}}_{t}$ of an LRNN layer in (1), we use the function $\mathrm{dec}$, to abstract away the computations that are done on ${\bm{H}}_{t}$ and ${\bm{x}}_{t}$, since they are not part of the recurrence. In this work, we do not consider the internal structure of $\mathrm{dec}$, but it usually contains a normalization and a feed-forward neural network and it can approximate any function.
@@ Appendix B Parity and Modular Counting – Proofs
Note that $\mathcal{D}_{\min}$ might not be a singleton. We naturally extend this function on complex numbers by applying it separately to the real and imaginary part, and then to complex-valued matrices by applying it element-wise. The following lemma is a key element of the proofs of Theorems 1 and 2. The sequence $a_{k}$ in the lemma takes the form of the imaginary or real part of the elements of the $k$-th power of a matrix with real eigenvalues ($\lambda_{i}$ will be one eigenvalue), expressed using the Jordan canonical form. See Section B.1 for more details on the Jordan Canonical Form. Intuitively, the lemma shows that if some of the $\lambda_{i}$-s are negative then for $k$ large enough, $a_{k}$ in finite precision will alternate between two values. Instead, if the $\lambda_{i}$-s are only nonnegative, $a_{k}$ in finite precision becomes constant for large enough $k$.
@@ B.1 Proof of Theorem 1
The language $(11)^{*}$ contains all sequences with an even number of ones. For the sequence of all ones $1^{k}$, an FSA recognizing the language, for the sequence $1^{k}$ will output $y_{k}=1$ if $k$ is even and $y_{k}=0$ if $k$ is odd. Consider an LRNN with one layer as in (1). We will prove that if ${\bm{A}}(1)$ has only nonnegative eigenvalues, then there exists a $\overline{k}>0$ such that for every $k\geq\overline{k}$, the finite precision version of the state ${\bm{H}}_{k}$ corresponding to the sequence $1^{k}$ does not depend on $k$ and is equal to $\overline{{\bm{H}}}$. Hence, no matter the choice of $\mathrm{dec}$, also the finite precision version of $\hat{{\bm{y}}}_{k}$ will not vary with $k$ and thus for some $k^{\prime}\geq\bar{k}$, $\hat{\bm{y}}_{k^{\prime}}\neq k^{\prime}\mod 2=y_{k^{\prime}}$. An inductive argument can then be used for the case of LRNNs with multiple (finitely many) layers, using the fact that the input of the next layer will be constant for $k$ large enough, as the input of the first layers.
@@ B.2 Proof of Theorem 2
Therefore, for $k\geq\bar{k}_{2}$, the function $k\mapsto\overline{\bm{H}}^{(2)}_{k}$ will be periodic with period a divisor of four and hence no matter the choice of $\mathrm{dec}^{(2)}$, also the function $k\mapsto\hat{\bm{y}}^{(2)}_{k}$ will be periodic with period a divisor of $4$. Consequently, with two layers one can recognize the language $(1^{m})^{*}$ only when $m=1$, $m=2$, or $m=4$, since those are the only cases where $k\mapsto y_{k}$ has a period which is a divisor of $4$. Thanks to the assumption on the eigenvalues of the products of state-transition matrices, we can extend this argument inductively to the case of an LRNN with $L$ layers. In particular, for the $i$-th layer, the induction hypothesis is that we assume $k\mapsto{\bm{x}}_{k}^{(i)}$, mapping $k$ to the $k$-th input to the layer, to be periodic with period a divisor of $2^{i-1}$ for $k$ large enough. Hence, there will be $2^{i-1}$ subsequences of states containing powers of the product of $2^{i-1}$ state-transition matrices. From our hypothesis on the eigenvalues of products of state-transition matrices, such product will have only real eigenvalues and hence each subsequence will have 2 converging subsequences resulting in $k\mapsto{\bm{H}}_{k}^{(i)}$ and consequently $k\mapsto\hat{\bm{y}}_{k}^{(i)}$ and hence $k\mapsto{\bm{x}}_{k}^{(i+1)}$, for $k$ large enough, being periodic with period a divisor of $2^{i}$. Therefore, for the $L$-th layer, there exists $\bar{k}_{L}\geq 0$ such that for every $k\geq\bar{k}_{L}$, the function $k\mapsto\hat{\bm{y}}^{(L)}_{k}$ is periodic with a period which is a divisor of $2^{L}$ and thus it can recognize the language $(1^{m})^{*}$ only when $02^{L}\mod m=$ 0, which happens only when there exists $p\leq L$ such that $m=2^{p}$ and hence $m$ is a power of two, ending the proof. ∎
@@ C.1 Proof of Proposition 1
where ${\bm{e}}_{i}$ is the $i$-th element of the canonical basis of $\mathbb{R}^{n}$. Hence, ${\bm{S}}\in\mathcal{M}_{n}^{n}([0,1])$. The proof of the first part is concluded since we wrote each of ${\bm{U}},{\bm{S}},{\bm{V}}$ as a product of at most $n$ GH matrices. If ${\bm{M}}$ is orthogonal we apply the $n$-reflections theorem directly. We also note that if ${\bm{M}}={\bm{P}}\in\{0,1\}^{n\times n}$ with ${\bm{P}}$ being a permutation matrix different from the identity, it can be written as products of at most $n-1$ swaps, i.e. permutation matrices permuting only two elements. Therefore we have that there exists an integer $k\leq n-1$ and indices $i_{1},\dots,i_{k}$ and $j_{1},\dots,j_{k}$ such that $i_{l}\neq j_{l}$ and
@@
Third item Let ${\bm{N}}={\bm{C}}_{1}{\bm{C}}_{2}\cdots{\bm{C}}_{k}\in\mathcal{M}_{k}^{n}((-1,1])$, with ${\bm{C}}_{i}={\bm{I}}-\beta_{i}{\bm{z}}_{i}{\bm{z}}_{i}^{\top}$ with $\left\lVert{\bm{z}}_{i}\right\rVert=1$ and $\beta_{i}\in[0,2)$. If ${\bm{N}}={\bm{I}}$ the statement is satisfied, otherwise, let $\mathcal{V}=\mathrm{span}\{{\bm{z}}_{i}\,:\,i\in\{1,\dots,k\},\beta_{i}>0\}$. Any unit vector ${\bm{v}}\in\mathbb{R}^{n}$ can then be written as ${\bm{v}}={\bm{v}}_{1}+{\bm{v}}_{2}$ with ${\bm{v}}_{1}\in\mathcal{V}$, ${\bm{v}}_{2}\in\mathcal{V}^{\top}$ and $\left\lVert{\bm{v}}_{1}\right\rVert,\left\lVert{\bm{v}}_{2}\right\rVert\leq 1$. Now, if ${\bm{v}}_{1}=0$, then ${\bm{N}}{\bm{v}}={\bm{v}}$, and hence ${\bm{v}}$ is an eigenvector with eigenvalue $1$. Instead, if ${\bm{v}}_{1}\neq 0$, then there exists $i^{\prime}\in\{1,\dots,k\}$ (we take the largest one one) such that $\beta_{i^{\prime}}\in(0,2)$ and ${\bm{v}}^{\top}{\bm{z}}_{i^{\prime}}={\bm{v}}_{1}^{\top}{\bm{z}}_{i^{\prime}}\in(0,1]$ and if $i^{\prime}<k$, then either $\beta_{j}=0$ or ${\bm{z}}_{j}^{\top}{\bm{v}}=0$ so that ${\bm{C}}_{j}{\bm{v}}={\bm{v}}$ for all $j\in\{i^{\prime}+1,\dots,k\}$. Moreover, we have that
@@
where the last line comes from the fact that $\min_{x\in[0,2]}x(2-x)=1$ and is only reached at $x=0$ and $x=2$, while $\beta_{i^{\prime}}\in(0,2)$. Therefore, since for every $i$, $\left\lVert{\bm{C}}_{i}\right\rVert\leq 1$ and the Euclidean norm is sub-multiplicative we have
@@ C.2 Proof of Theorem 3
We view the LRNN layer in (1) as the automaton $\mathcal{A}_{\mathrm{lin}}=(\Sigma,\mathcal{H},{\bm{H}}_{0},\delta_{\mathrm{lin}})$, where $\delta_{\mathrm{lin}}({\bm{H}},w)={\bm{A}}(w){\bm{H}}+{\bm{B}}(w)$, which is extended in the usual way, and =$\mathcal{H}=\{\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Since we assumed that $\mathcal{T}(\mathcal{A})$ is a group, from Cayley’s theorem we have that it is isomorphic to a subgroup of $S_{n}$, which is the set of permutations on a set of $n$ elements. Furthermore, each element in $S_{n}$ can be represented as an $n\times n$ permutation matrix. Since in general $n\neq|Q|$, we cannot let $\mathcal{H}$ to be a set of one hot vectors each corresponding to states in $Q$. Instead, we let ${\bm{H}}_{0}=(1,\dots,n)^{\top}$, $\mathcal{P}\subset\{0,1\}^{n\times n}$ be the set of permutation matrices and set ${\bm{B}}\equiv 0$ and ${\bm{A}}:\Sigma\to\mathcal{P}$ to be the function mapping each letter $w\in\Sigma$ to the permutation matrix corresponding to $\delta(\cdot,w)$. With this choice we can see that the function $f:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{T}(\mathcal{A})$ such that $f(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta(\cdot,{\bm{w}})$ for every ${\bm{w}}\in\Sigma^{*}$ is one-to-one (biejctive), and from our choice of ${\bm{H}}_{0}$, the map $h:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{H}$ such that for every ${\bm{w}}\in\Sigma^{*}$, $h(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})$ is also bijective. Moreover, the map $\phi:\mathcal{T}(\mathcal{A})\to Q$ such that $\phi(\delta(\cdot,{\bm{w}}))=\delta(q_{0},{\bm{w}})$ is surjective because we consider states that are only reachable from the initial state $q_{0}$, i.e. $Q=\{\delta(q_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Hence if we set $g=\phi\circ f\circ h^{-1}$, then $g:\mathcal{H}\to Q$ is surjective and for every $w\in\Sigma$ and ${\bm{H}}\in\mathcal{H}$ we have that

### v3

@@ Acknowledgments
We would like to thank David Salinas, Herilalaina Rakotoarison, Eric Alcaide, Arya Akhavan, Matia Bojovic, Erfan Mirzaei and the active members of the Flash Linear Attention discord channel for their constructive discussions and feedback. We acknowledge the support and assistance of the Data Science and Computation Facility and its Support Team, in particular Mattia Pini, in utilizing the IIT High-Performance Computing Infrastructure, on which we run our largest experiments. This research was partially supported by the following sources: PNRR MUR Project PE000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)“, funded by the European Union – NextGenerationEU, and EU Project ELSA under grant agreement No. 101070617. TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Frank Hutter acknowledges financial support by the Hector Foundation. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. The authors gratefully acknowledge the Gauss Center for Supercomputing eV (www.gauss-centre.eu) for funding this project by providing computing time on the GCS supercomputer JUWELS at Jülich Supercomputing Center (JSC). The MATH-HARD dataset which we use in one of our experiments was compiled from AoPS & the AoPS Community, MATHCOUNTS, the MAA, the Centre for Education in Mathematics and Computing, the Harvard-MIT Math Tournament, the Math Prize for Girls, MOEMS, the Mandelbrot Competition, and the Institute of Mathematics and Applications. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
@@
  • Appendix A contains additional details on the notation used, on Table 1, on the relationship between RNNs and regular languages, on the assumption of finite precision, on the states, and on the function $\mathrm{dec}$.
@@ A.2 Details of Table 1
The Mamba recurrence in Equations 3 and 4 in (Gu & Dao, 2024) is applied independently to each channel of the input sequence. Expressing the full recurrence in the matrix-form of (1) is challenging, as it would require concatenating the rows of the matrix ${\bm{H}}_{t}$. For simplicity, in Table 1 we write instead the recurrence for each row of ${\bm{H}}_{t}$. In particular, Let ${\bm{x}}_{t}\in\mathbb{R}^{d}$ be the input of the layer, ${\bm{W}}_{\Delta}\in\mathbb{R}^{d\times d}$, ${\bm{w}}_{2}\in\mathbb{R}^{d}$, ${\bm{W}}_{1}=({\bm{w}}_{1},\dots,{\bm{w}}_{n})^{\top}\in\mathbb{R}^{n\times d}$ be learnable parameters, ${\bm{q}}_{t}\in\mathbb{R}^{n},{\bm{k}}_{t}=(k_{t,1},\dots,k_{t,n})^{\top}\in\mathbb{R}^{n}$ be learnable functions of the input and $\bm{\Delta}_{t}=\mathrm{softplus}({\bm{W}}_{\Delta}{\bm{x}}_{t})$. Then, if we set ${\bm{H}}_{t}=({\bm{h}}_{t,1},\dots,{\bm{h}}_{t,n})^{\top}\in\mathbb{R}^{n\times d}$ and ${\bm{H}}_{0}=0$, we can write the recurrence for the $i$-th row of ${\bm{H}}_{t}$ and the output as
@@
where $\mathbf{1}$ is the vector of $n$ ones. However, such a recurrence is not in the form (1), since we have replaced the matrix-matrix product ${\bm{A}}({\bm{x}}_{t}){\bm{H}}_{t}$ with the element-wise product ${\bm{A}}({\bm{x}}_{t})\odot{\bm{H}}_{t}$. Note that we follow the implementation of ${\bm{B}}({\bm{x}}_{t})$ used in the official Mamba codebase, which simplifies the expression originally presented in Equation 4 of (Gu & Dao, 2024) as described by the authors in a GitHub Issue [^3] https://github.com/state-spaces/mamba/issues/19.
@@ A.3 Regular Languages and Recurrent Neural Networks
Clearly, any FSA can be implemented by an RNN layer if $g$ is sufficiently expressive to model its state transition function.
@@ A.4.1 Initial State, Matrix-valued States, and The Decoder function
While we mention that the states ${\bm{H}}_{t}$ are generally matrices of dimension $n\times d$, for our theoretical constructions (excluding the first two theorems), we set $d=1$, so that states are vector-valued. Hence, for the problems that we consider, we find that having a matrix-valued state ($d>1$) brings no theoretical advantage, while it is very important for associative recall.
To compute the output $\hat{\bm{y}}_{t}$ from the state ${\bm{H}}_{t}$ and the vector ${\bm{x}}_{t}$ of an LRNN layer in (1), we use the function $\mathrm{dec}$, to abstract away the computations that are done on ${\bm{H}}_{t}$ and ${\bm{x}}_{t}$, since they are not part of the recurrence. In this work, we do not consider the internal structure of $\mathrm{dec}$, but it usually contains a normalization and a feed-forward neural network and it can approximate any continuous function.
@@ Appendix B Parity and Modular Counting – Proofs
Note that $\mathcal{D}_{\min}$ might not be a singleton. We naturally extend this function on complex numbers by applying it separately to the real and imaginary part, and then to complex-valued matrices by applying it element-wise. The following lemma is a key element of the proofs of Theorems 1 and 2. There, the sequence $a_{k}$ in the lemma takes the form of the imaginary or real part of the elements of the $k$-th power of a matrix with real eigenvalues ($\lambda_{i}$ will be one eigenvalue), expressed using the Jordan canonical form. See Section B.1 for more details on the Jordan Canonical Form. Intuitively, the lemma shows that if some of the $\lambda_{i}$-s are negative then for $k$ large enough, $a_{k}$ in finite precision will alternate between two values. Instead, if the $\lambda_{i}$-s are only nonnegative, $a_{k}$ in finite precision becomes constant for large enough $k$.
@@ B.1 Proof of Theorem 1
The language $(11)^{*}$ contains all sequences with an even number of ones. An FSA recognizing the language, for the sequence $1^{k}$ will output $y_{k}=1$ if $k$ is even and $y_{k}=0$ if $k$ is odd. Consider an LRNN with one layer as in (1). We will prove that if ${\bm{A}}(1)$ has only nonnegative eigenvalues, then there exists a $\overline{k}>0$ such that for every $k\geq\overline{k}$, the finite precision version of the state ${\bm{H}}_{k}$ corresponding to the sequence $1^{k}$ does not depend on $k$ and is equal to $\overline{{\bm{H}}}$. Hence, no matter the choice of $\mathrm{dec}$, also the finite precision version of $\hat{{\bm{y}}}_{k}$ will not vary with $k$ and thus for some $k^{\prime}\geq\bar{k}$, $\hat{\bm{y}}_{k^{\prime}}\neq k^{\prime}\mod 2=y_{k^{\prime}}$. An inductive argument can then be used for the case of LRNNs with multiple (finitely many) layers, using the fact that the input of the next layer will be constant for $k$ large enough, as the input of the first layers.
@@ B.2 Proof of Theorem 2
Therefore, for $k\geq\bar{k}_{2}$, the function $k\mapsto\overline{\bm{H}}^{(2)}_{k}$ will be periodic with period a divisor of four and hence no matter the choice of $\mathrm{dec}^{(2)}$, also the function $k\mapsto\hat{\bm{y}}^{(2)}_{k}$ will be periodic with period a divisor of $4$. Consequently, with two layers one can recognize the language $(1^{m})^{*}$ only when $m=1$, $m=2$, or $m=4$, since those are the only cases where $k\mapsto y_{k}$ has a period which is a divisor of $4$. Thanks to the assumption on the eigenvalues of the products of state-transition matrices, we can extend this argument inductively to the case of an LRNN with $L$ layers. In particular, for the $i$-th layer, the induction hypothesis is that we assume $k\mapsto{\bm{x}}_{k}^{(i)}$, mapping $k$ to the $k$-th input to the layer, to be periodic with period a divisor of $2^{i-1}$ for $k$ large enough. Hence, there will be $2^{i-1}$ subsequences of states, each containing powers of the product of $2^{i-1}$ state-transition matrices. From our hypothesis on the eigenvalues of products of state-transition matrices, such product will have only real eigenvalues and hence each subsequence will have 2 converging subsequences resulting in $k\mapsto{\bm{H}}_{k}^{(i)}$ and consequently $k\mapsto\hat{\bm{y}}_{k}^{(i)}$ and hence $k\mapsto{\bm{x}}_{k}^{(i+1)}$, for $k$ large enough, being periodic with period a divisor of $2^{i}$. Therefore, for the $L$-th layer, there exists $\bar{k}_{L}\geq 0$ such that for every $k\geq\bar{k}_{L}$, the function $k\mapsto\hat{\bm{y}}^{(L)}_{k}$ is periodic with a period which is a divisor of $2^{L}$ and thus it can recognize the language $(1^{m})^{*}$ only when $02^{L}\mod m=$ 0, which happens only when there exists $p\leq L$ such that $m=2^{p}$ and hence $m$ is a power of two, ending the proof. ∎
@@ C.1 Proof of Proposition 1
where ${\bm{e}}_{i}$ is the $i$-th element of the canonical basis of $\mathbb{R}^{n}$. Hence, ${\bm{S}}\in\mathcal{M}_{n}^{n}([0,1])$. The proof of the first part is concluded since we wrote each of ${\bm{U}},{\bm{S}},{\bm{V}}$ as a product of at most $n$ GH matrices. If ${\bm{M}}$ is orthogonal, we apply the $n$-reflections theorem directly. We also note that if ${\bm{M}}={\bm{P}}\in\{0,1\}^{n\times n}$ with ${\bm{P}}$ being a permutation matrix different from the identity, it can be written as products of at most $n-1$ swaps, i.e. permutation matrices permuting only two elements. Therefore we have that there exists an integer $k\leq n-1$ and indices $i_{1},\dots,i_{k}$ and $j_{1},\dots,j_{k}$ such that $i_{l}\neq j_{l}$ and
@@
Third item Let ${\bm{N}}={\bm{C}}_{1}{\bm{C}}_{2}\cdots{\bm{C}}_{k}\in\mathcal{M}_{k}^{n}((-1,1])$, with ${\bm{C}}_{i}={\bm{I}}-\beta_{i}{\bm{z}}_{i}{\bm{z}}_{i}^{\top}$ with $\left\lVert{\bm{z}}_{i}\right\rVert=1$ and $\beta_{i}\in[0,2)$. If ${\bm{N}}={\bm{I}}$ the statement is satisfied, otherwise, let $\mathcal{V}=\mathrm{span}\{{\bm{z}}_{i}\,:\,i\in\{1,\dots,k\},\beta_{i}>0\}$. Any unit vector ${\bm{v}}\in\mathbb{R}^{n}$ can then be written as ${\bm{v}}={\bm{v}}_{1}+{\bm{v}}_{2}$ with ${\bm{v}}_{1}\in\mathcal{V}$, ${\bm{v}}_{2}\in\mathcal{V}^{\top}$ and $\left\lVert{\bm{v}}_{1}\right\rVert,\left\lVert{\bm{v}}_{2}\right\rVert\leq 1$. Now, if ${\bm{v}}_{1}=0$, then ${\bm{N}}{\bm{v}}={\bm{v}}$, and hence ${\bm{v}}$ is an eigenvector with eigenvalue $1$. Instead, if ${\bm{v}}_{1}\neq 0$, then there exists $i^{\prime}\in\{1,\dots,k\}$ (we take the largest one) such that $\beta_{i^{\prime}}\in(0,2)$ and $({\bm{v}}^{\top}{\bm{z}}_{i^{\prime}})^{2}=({\bm{v}}_{1}^{\top}{\bm{z}}_{i^{\prime}})^{2}\in(0,1]$. Therefore, if $i^{\prime}<k$, then either $\beta_{j}=0$ or ${\bm{z}}_{j}^{\top}{\bm{v}}=0$ so that ${\bm{C}}_{j}{\bm{v}}={\bm{v}}$ for all $j\in\{i^{\prime}+1,\dots,k\}$. Moreover, we have that
@@
where the last line comes from the fact that $\min_{x\in[0,2]}x(2-x)=0$ and is only reached at $x=0$ and $x=2$, while $\beta_{i^{\prime}}\in(0,2)$. Therefore, since for every $i$, $\left\lVert{\bm{C}}_{i}\right\rVert\leq 1$ and the Euclidean norm is sub-multiplicative we have
@@ C.2 Proof of Theorem 3
We view the LRNN layer in (1) as the automaton $\mathcal{A}_{\mathrm{lin}}=(\Sigma,\mathcal{H},{\bm{H}}_{0},\delta_{\mathrm{lin}})$, where $\delta_{\mathrm{lin}}({\bm{H}},w)={\bm{A}}(w){\bm{H}}+{\bm{B}}(w)$, which is extended in the usual way, and =$\mathcal{H}=\{\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Since we assumed that $\mathcal{T}(\mathcal{A})$ is a group, from Cayley’s theorem we have that it is isomorphic to a subgroup of $S_{n}$, which is the set of permutations on a set of $n$ elements. Furthermore, each element in $S_{n}$ can be represented as an $n\times n$ permutation matrix. Since in general $n\neq|Q|$, we cannot let $\mathcal{H}$ to be a set of one hot vectors each corresponding to states in $Q$. Instead, we let ${\bm{H}}_{0}=(1,\dots,n)^{\top}$, $\mathcal{P}\subset\{0,1\}^{n\times n}$ be the set of permutation matrices and set ${\bm{B}}\equiv 0$ and ${\bm{A}}:\Sigma\to\mathcal{P}$ to be the function mapping each letter $w\in\Sigma$ to the permutation matrix corresponding to $\delta(\cdot,w)$. With this choice we can see that the function $f:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{T}(\mathcal{A})$ such that $f(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta(\cdot,{\bm{w}})$ for every ${\bm{w}}\in\Sigma^{*}$ is one-to-one (bijective), and from our choice of ${\bm{H}}_{0}$, the map $h:\mathcal{T}(\mathcal{A}_{\mathrm{lin}})\to\mathcal{H}$ such that for every ${\bm{w}}\in\Sigma^{*}$, $h(\delta_{\mathrm{lin}}(\cdot,{\bm{w}}))=\delta_{\mathrm{lin}}({\bm{H}}_{0},{\bm{w}})$ is also bijective. Moreover, the map $\phi:\mathcal{T}(\mathcal{A})\to Q$ such that $\phi(\delta(\cdot,{\bm{w}}))=\delta(q_{0},{\bm{w}})$ is surjective because without loss of generality we can consider states that are only reachable from the initial state $q_{0}$, i.e. $Q=\{\delta(q_{0},{\bm{w}})\,:\,{\bm{w}}\in\Sigma^{*}\}$. Hence if we set $g=\phi\circ f\circ h^{-1}$, then $g:\mathcal{H}\to Q$ is surjective and for every $w\in\Sigma$ and ${\bm{H}}\in\mathcal{H}$ we have that